# More Likely Than You Think: Inclination-Driving Secular Resonances are Common in Known Exoplanet Systems

### v1

@@ Abstract
Multi-planet systems face significant challenges to detection. For example, further orbiting planets have reduced signal-to-noise ratio in radial velocity detection methods, and small mutual inclinations between planets can prevent them from all transiting. One mechanism to excite mutual inclination between planets is secular resonance, where the nodal precession frequencies of the planets align such as to greatly increase the efficiency of angular momentum transport between planets. These resonances can significantly misalign planets from one another, hindering detection, and typically can only occur when there are three or more planets in the system. Naively, systems can only be in resonance for particular combinations of planet semimajor axes and masses; however, effects that alter the nodal precession frequencies of the planets, such as the decay of stellar oblateness, can significantly expand the region of parameter space where resonances occur. In this work, we explore known three-planet systems, determine whether they are in (or were in) secular resonance due to evolving stellar oblateness, and demonstrate the implications of resonance on their detectability and stability. We show that about $20\%$ of a sample of three planet transiting systems seem to undergo sweeping resonances early in their lives.
@@ 1 Introduction
$\dot{\Omega}_{J_{2}},\,\,\dot{\omega}_{J_{2}}=\mp n\left[\frac{3}{2}J_{2}\left(\frac{R_{\mathrm{\star}}}{a}\right)^{2}\right]\propto\mp J_{2}P^{-7/3},$ (1)
@@ 2 Secular Resonances and the Relation to $J_{2}$
In Figure 3, we show the time evolution of Kepler-9’s planetary system. Specifically, we show the mutual inclination between the planets (left top panel) and eccentricity (left bottom panel) as well as the alignment or misalignment of the outer two planets’ longitude of ascending notes (top right panel) and longitudes of periapsis (bottom right panel). Mutual inclinations were obtained by combining $I_{j}$ and $\Omega_{j}$ to get the 3D planetary angular momentum vector for each planet and the mutual inclination was calculated using a dot product. The initial mutual inclination between the outer planets is chosen to be $2.5^{\circ}$. For simplicity, we adopt that the initial longitudes of ascending node of the planets are set to (approximately) zero. This evolution was obtained by integrating Equations 11 (which include $J_{2}$ precession) and adding GR precession. As Equations 11 are secularly averaged, the evolution in Figure 3 is therefore purely secular.
@@
Overplotted on all figures are the precession timescales for $J_{2}$ precession ($\tau_{J_{2}}$) on the inner planet (dashed red lines), and Laplace-Lagrange (LL) precession on the inner planet ($\tau_{LL}$) (approximated by $A_{00}(=-B_{00}$), solid red lines).
@@
Initially, $J_{2}$ is high, and the eccentricities and mutual inclinations oscillate about their initial value and do not evolve significantly for the first few Myr. This is because the $J_{2}$ evolution in Figure 2 remains constant for the first 1010\, Myr111see Appendix B for a discussion. Then, Kepler-9 is swept into a resonance, exciting Kepler-9d’s eccentricity and inclination, and the outer two planets’ longitudes of ascending nodes lock together ($\Omega_{b}=\Omega_{c}$). A similar effect occurs in the eccentricities and longitudes of pericenter ($\varpi_{b}=\varpi_{c}$), though at a later time in the evolution. The inclination excitation occurs at $t\approx 17\,$Myr and the eccentricity excitation at  $\approx\,22$ Myr.
@@ 3 The Effect of Initial Conditions on Secular Resonances
In the leftmost region, $a_{2}>>a_{1}$, there is no resonance, and inclinations induced on the inner planet are small. To the right of this region is a thin strip where inner planet inclination is maximized. This occurs for only a narrow subset of choices of $a_{1}$ and $a_{2}$. This region is caused by secular resonances when $J_{2}=0$ or when $J_{2}$ is sufficiently small to be negligible. To the right of the $J_{2}=0$ resonant region is a much larger region of enhanced inclination. This is caused by secular resonances for a range of nonzero, nontrivial $J_{2}$ values. This region is not in resonance for $J_{2}=0$, but for every point in the region, there exists some $J_{2}\neq 0$ which precesses the planets just enough to bring the planets into secular resonance. We call this the “Sweeping Resonance” region because for a particular value of $J_{2}$, only a small portion of the region is in resonance, and it is only as $J_{2}$ continuously evolves that that resonant portion “sweeps” across the parameter space. To the right of this region is a nonresonant region where $a_{1}\approx a_{2}$. To the right of that region, indicated by the dark lower triangle, is a forbidden region by the restriction that $a_{2}>a_{1}$. Written in the left panel are descriptions of how these regions’ boundaries change for different parameter values. For increasing $a_{0}$, the sweeping resonant region for $J_{2}>0$ decreases in size, but the $J_{2}=0$ resonance stays in place. For increasing maximum $J_{2}$, $M_{\star}$, and $R_{\star}$, the sweeping resonant region for $J_{2}>0$ increases in size.
@@ 4 Observed Systems
To quantify which systems undergo systematic misalignments similar to those depicted in Figure 3, for each of these 31 systems, we run 100 integrations of the secular equations described in Eqs 11. In these integrations, we include the effects of precession due to $J_{2}$ and GR, and we let $J_{2}$ evolve according to the upper curve of Figure 2. Each simulation was allowed to run for 22\,Gyr. In each simulation the initial conditions of the planets were randomized according to the prescription given in Table 2. We find that six of these 31 systems exhibit permanent misalignments between planets mediated by $J_{2}$. Similar to the evolution depicted in Figure 3, when $J_{2}$ passes a critical value, the outer two planets in these systems align their orbits. Simultaneously, the innermost planet becomes misaligned from the outer two. The values of $J_{2}$ and the corresponding times when the misalignment occurs are listed in Table 3.
@@
In Figure 6, we show the structure in $a_{1}/a_{0}$$a_{2}/a_{0}$ space of nine systems. Each panel of Figure 6 represents a different observed system, with the observed values of the semimajor axis ratios (derived from the extremely accurately measured period ratios) marked with a red dot. The top two rows of Figure 6 are the six systems from our $31$-system sample which were found in the Monte Carlo simulations to undergo permanent misalignments. The bottom row of Figure 6 contains three of the $25$ systems whose simulations did not undergo permanent misalignments. The bottom row systems are chosen from various regions of non-sweeping parameter space: high $a_{1}/a_{0}$ (Kepler-616; see Table 5), “static” resonance that does not sweep, and therefore does not produce misalignments (TOI-2076; discussed in detail in Section 4.2), and low $a_{1}/a_{0}$ (HD 22946; see Table 5).
@@
In each panel of Figure 7, we show the relationship between the initial mutual inclination between the outer two planets (x-axis) and the change in planets’ relative alignment (y-axes, $\Delta i_{01}^{mut}$ in red and $\Delta i_{12}^{mut}$ in blue). The top two rows are systems whose positions in $a_{1}/a_{0}$$a_{2}/a_{0}$ space indicate they undergo sweeping resonances. For these systems, there is an approximately linear relationship between initial mutual inclination between outer planets, and the (mis-)alignments between the planets after $J_{2}$ evolution has had its effect. In each panel, the dotted red line represents the minimum mutual inclination between the middle and inner planets such that the inner planet may not co-transit with the middle planet (under the simplifying assumption that the middle planet’s transit impact parameter $b=0$). Many of the sweeping resonant systems would require very large initial mutual inclinations between the outer two planets to make the inner planet no longer transit.
[…]
@@
Figure 7: Relation between initial mutual inclination between the two outermost planets and changes in alignment between planets over the course of the simulation for nine systems. The top six systems undergo resonance-mediated misalignments. The bottom three do not. In red circles are, for each simulation, the change in alignment between the inner and middle planets–the average mutual inclination at late times between the inner and middle planet minus the initial mutual inclination between the inner and middle planet. The red dotted line represents the minimum $\Delta i_{mut,01}$ such that the inner planet fails to transit at the end of the simulation (under the simplifying assumption that the middle planet has a transit impact parameter $b=0$). In blue triangles are, for each simulation, the change in alignment between the middle and outer planets–the average mutual inclination at late times between the middle and outer planet minus the initial mutual inclination between the middle and outer planet. The black dotted line represents the maximum value of $\Delta i_{mut,12}$ as a function of $i_{mut,12,init}$ such that the outer two planets both transit at the end of the simulation. The scheme by which orbital parameters were varied between simulations is described in Table 2.
@@ 4.2 The Impact of Static vs Sweeping Resonances
Static resonances notably do not produce the misalignment from evolving $J_{2}$ that sweeping resonances do. For example, the system TOI-2076 (observed by e.g., Osborn et al., 2022; Hedges et al., 2021) lies on such a static resonance as depicted in the bottom middle panel of Figure 6 . In Figure 9, we show the semimajor axis parameter space of TOI-2076 for $J_{2}=10^{-3}$ (left panel) and $J_{2}=0$ (right panel). The location of TOI-2076’s observed semimajor axis ratios is marked with a red dot. In both panels, TOI-2076 lies on a secular resonance with the same expected inner planet maximum inclination. This indicates that there is no expected change to the system as $J_{2}$ evolves from $\sim 10^{-3}$ to 0. Therefore, we do not expect the inner planet to systematically and/or permanently misalign from its companions any more than is provided in its initial conditions.
Indeed, in Figure 8, we show the time evolution of one of the simulations of TOI-2076 from the Monte Carlo. In the bottom left panel, we can see that the longitudes of ascending nodes between the outer two planets are locked together, indicating the resonance is active; however, the upper left panel shows that mutual inclinations do not significantly change. The planets do not misalign. In the bottom middle panel of Figure 7, we can see that even for large initial values of $i_{mut,12}$, no permanent misalignment occurs between the planets in this system–they retain whatever mutual inclinations they are initialized with.
@@
Overplotted on all figures are the precession timescales for $J_{2}$ precession on the inner planet, and Laplace-Lagrange (LL) precession on the inner planet (approximated by $A_{00}=-B_{00}$).

### v2

@@ Abstract
Multi-planet systems face significant challenges to detection. For example, further-orbiting planets have reduced signal-to-noise ratio in radial velocity detection methods, and small mutual inclinations between planets can prevent them from all transiting. One mechanism to excite mutual inclination between planets is secular resonance, where the nodal precession frequencies of the planets align such as to greatly increase the efficiency of angular momentum transport between planets. These resonances can significantly misalign planets from one another, hindering detection, and typically can only occur when there are three or more planets in the system. Naively, systems can only be in resonance for particular combinations of planet semimajor axes and masses; however, effects that alter the nodal precession frequencies of the planets, such as the decay of stellar oblateness, can significantly expand the region of parameter space where resonances occur. In this work, we explore known three-planet systems, determine whether they are in (or were in) secular resonance due to evolving stellar oblateness, and demonstrate the implications of resonance on their detectability and stability. We show that about $20\%$ of a sample of three planet transiting systems seem to undergo these resonances early in their lives.
@@ 1 Introduction
$\dot{\Omega}_{J_{2}},\,\,\dot{\varpi}_{J_{2}}=\mp n\left[\frac{3}{2}J_{2}\left(\frac{R_{\mathrm{\star}}}{a}\right)^{2}\right]\propto\mp J_{2}P^{-7/3},$ (1)
@@ 2 Secular Resonances and the Relation to $J_{2}$
In Figure 3, we show the time evolution of Kepler-9’s planetary system. Specifically, we show the mutual inclination between the planets (left top panel) and eccentricity (left bottom panel) as well as the alignment or misalignment of the outer two planets’ longitude of ascending notes (top right panel) and longitudes of periapsis (bottom right panel). Overplotted on all panels are the eigenfrequencies of the system (blue and red). Mutual inclinations were obtained by combining $I_{j}$ and $\Omega_{j}$ to get the 3D planetary angular momentum vector for each planet and the mutual inclination was calculated using a dot product. The initial mutual inclination between the outer planets is chosen to be $2.5^{\circ}$. For simplicity, we adopt that the initial longitudes of ascending node of the planets are set to (approximately) zero. This evolution was obtained by integrating Equations 11 (which include $J_{2}$ precession) and adding GR precession. As Equations 11 are secularly averaged, the evolution in Figure 3 is therefore purely secular.
@@
Overplotted on all figures are the Laplace-Lagrange eigenfrequencies $f_{i}$ and $g_{i}$.
@@
Initially, $J_{2}$ is high, and the eccentricities and mutual inclinations oscillate about their initial value and do not evolve significantly for the first few Myr. This is because the $J_{2}$ evolution in Figure 2 remains constant for the first 1010\, Myr.111see Appendix B for a discussion. The inner planet initially has the majority of its eccentricity and inclination oscillation amplitude in the shortest-period eigenmode. The outer two planets have vanishingly small amplitude in this mode. However, as $J_{2}$ evolves, for both inclination and eccentricity, this mode eventually has the same period as one of the other two modes–bringing the inner planet into resonance with its companions. Kepler-9 is swept into a resonance, exciting Kepler-9d’s eccentricity and inclination, and the outer two planets’ longitudes of ascending nodes lock together ($\Omega_{b}=\Omega_{c}$). A similar effect occurs in the eccentricities and longitudes of pericenter ($\varpi_{b}=\varpi_{c}$), though at a later time in the evolution. The inclination excitation occurs at $t\approx 17\,$Myr and the eccentricity excitation at  $\approx\,22$ Myr. Note that despite appearing instantaneous, the exchange of angular momentum that excites the inner planet takes place over $\sim 2\times 10^{5}$ years. However, despite the exchange occurring over such a long timescale, the exchange is not an adiabatic resonant capture–the inner planet does not remain in resonance with the outer two planets as the colliding eigenvalues diverge. The amplitude of the largest-amplitude eigenmode of the inner planet (i.e., $\max(e_{00},e_{01},e_{02})$ and $\max(I_{00},I_{01},I_{02})$ in Eq 16) is an adiabatic invariant before and after the resonant passage (see e.g., Ward et al., 1976, for a complete treatment), but transitions to a new value during the resonant passage. Moreover, the resonance is only ever achieved by the inner planet’s highest-amplitude eigenfrequency approaching another eigenfrequency from above (see the overplotted eigenmode periods in Figure 3). To approach from below would require $J_{2}$ to increase with time.
@@ 3 The Effect of Initial Conditions on Secular Resonances
In the leftmost region, $a_{2}>>a_{1}$, there is no resonance, and inclinations induced on the inner planet are small. To the right of this region is a thin strip where inner planet inclination is maximized. This occurs for only a narrow subset of choices of $a_{1}$ and $a_{2}$. This region is caused by secular resonances when $J_{2}=0$ or when $J_{2}$ is sufficiently small to be negligible. These $J_{2}=0$ resonances occur for the inner planet when the inner planet’s oscillation has majority amplitude in the same eigenmode as the outer planets at $J_{2}=0$, causing it to precess together with one or more of the outer planets. This increases the amount of eccentricity and inclination oscillation of the inner planet relative to the amount of oscillation if the semimajor axis of the inner planet was perturbed slightly (see e.g., Faridani et al., 2022). To the right of the $J_{2}=0$ resonant region is a much larger region of enhanced inclination. This is caused by secular resonances for a range of nonzero, nontrivial $J_{2}$ values. This region is not in resonance for $J_{2}=0$, but for every point in the region, there exists some $J_{2}\neq 0$ which precesses the planets just enough to bring the planets into secular resonance. We call this the “Sweeping Resonance” region because for a particular value of $J_{2}$, only a small portion of the region is in resonance, and it is only as $J_{2}$ continuously evolves that that resonant portion “sweeps” across the parameter space. To the right of this region is a nonresonant region where $a_{1}\approx a_{2}$. To the right of that region, indicated by the dark lower triangle, is a forbidden region by the restriction that $a_{2}>a_{1}$. Written in the left panel are descriptions of how these regions’ boundaries change for different parameter values. For increasing $a_{0}$, the sweeping resonant region for $J_{2}>0$ decreases in size, but the $J_{2}=0$ resonance stays in place. For increasing maximum $J_{2}$, $M_{\star}$, and $R_{\star}$, the sweeping resonant region for $J_{2}>0$ increases in size.
@@ 4 Observed Systems
To quantify which systems undergo systematic misalignments similar to those depicted in Figure 3, for each of these 31 systems, we run 100 integrations of the secular equations described in Eqs 11. In these integrations, we include the effects of precession due to $J_{2}$ and GR, and we let $J_{2}$ evolve according to the upper curve of Figure 2. Each simulation was allowed to run for 22\,Gyr. In each simulation the initial conditions of the planets were randomized according to the prescription given in Table 2. We find that six of these 31 systems exhibit permanent misalignments between planets mediated by $J_{2}$, corresponding to about one fifth of the sample. Similar to the evolution depicted in Figure 3, when $J_{2}$ passes a critical value, the outer two planets in these systems align their orbits. Simultaneously, the innermost planet becomes misaligned from the outer two. The values of $J_{2}$ and the corresponding times when the misalignment occurs are listed in Table 3.
@@
In Figure 6, we show the structure of increases of inner-planet inclination in $a_{1}/a_{0}$$a_{2}/a_{0}$ space of nine systems. Each panel of Figure 6 represents a different observed system, with the observed values of the semimajor axis ratios (derived from the extremely accurately measured period ratios) marked with a red dot. The top two rows of Figure 6 are the six systems from our $31$-system sample which were found in the Monte Carlo simulations to undergo permanent misalignments. The bottom row of Figure 6 contains three of the $25$ systems whose simulations did not undergo permanent misalignments. The bottom row systems are chosen from various regions of non-sweeping parameter space: high $a_{1}/a_{0}$ (Kepler-616; see Table 5), “static” resonance that does not sweep, and therefore does not produce misalignments (TOI-2076; discussed in detail in Section 4.2), and low $a_{1}/a_{0}$ (HD 22946; see Table 5). The values of $i_{0,max}$ shown in Figure 6 are calculated using the method described at the end of Section 2 for initial mutual inclinations between middle and outer planets of $3^{\circ}$. The predictions of inner-planet inclination of Figure 6 and the simulations shown later in the red circles of Figure 7 demonstrate good agreement between the analytic prediction and the simulated result.
@@
In each panel of Figure 7, we show the relationship between the initial mutual inclination between the outer two planets (x-axis) and the change in planets’ relative alignment (y-axes, $\Delta i_{01}^{mut}$ in red and $\Delta i_{12}^{mut}$ in blue). The top two rows are systems whose positions in $a_{1}/a_{0}$$a_{2}/a_{0}$ space indicate they undergo sweeping resonances. For these systems, there is an approximately linear relationship between initial mutual inclination between outer planets, and the (mis-)alignments between the planets after $J_{2}$ evolution has had its effect. In each panel, the dotted red line represents the minimum mutual inclination between the middle and inner planets such that the inner planet may not co-transit with the middle planet under the simplifying assumption that the middle planet’s transit impact parameter $b_{1}=0$. Many of the sweeping resonant systems would require very large initial mutual inclinations between the outer two planets to make the inner planet no longer transit.
However, that is not the only constraint we can place on the initial conditions. If the outer two planets are prepared with high initial mutual inclinations, under what conditions do they align enough over the course of the simulation such that at the end, they both transit? This is shown in the blue dotted lines. For blue triangles above the blue dotted line, the initial mutual inclination between the outer two planets is high enough that they may not transit together (assuming the middle planet’s impact parameter $b_{1}=0$), even after the sweeping resonance from evolving $J_{2}$ partially aligns their orbits. Of course, if the longitudes of ascending nodes are favorable, even two greatly mutually inclined planets may still co-transit. Therefore, even systems with red dots above the red dotted line or blue triangles above the blue dotted line may have all three planets transit depending on nodal positioning. To account for these two constraints on $i_{mut,12,init}$, regions on Figure 7 are shaded. The shaded regions correspond to the ranges of $i_{mut,12,init}$ such that all three planets are guaranteed to co-transit under the $b_{1}=0$ assumption. The color of the shaded region indicates which planet escapes transit first as $i_{mut,12,init}$ increases–red for the inner planet and blue for the outer planet. Therefore, to be guaranteed consistency with observations, $i_{mut,12,init}$ must be smaller than the rightmost edge of the shaded region. For example, Kepler-487 (which experiences sweeping resonance), has only a small range of $i_{mut,12,init}$ values where co-transit is guaranteed. The outer planets of Kepler-487 do not align sufficiently during resonance to overcome large initial mutual inclinations, and at $i_{mut,12,init}\sim 2^{\circ}$ the outer two planets no longer always co-transit. Systems that do not undergo sweeping resonance–such as Kepler-616, for example–have far stricter restraints on $i_{mut,12,init}$, as they are subject to the geometric requirements for co-transit without a sweeping resonance to align the outer planets. Additionally, Kepler-487, as well as K2-198 and EPIC 249893012, also has an interesting property where if $i_{mut,12,init}$ is near zero, then the inner and middle planets will be the ones aligning, and the outer planet misaligns from the rest. This is represented by negative (positive) values of $\Delta i_{01}^{mut}$ ($\Delta i_{12}^{mut}$).
With these constraints on initial inclinations in mind, it may be worthwhile to consider whether they are consistent with known constraints on the exoplanet population. For example, He et al. (2020) and Millholland et al. (2021) find that the Kepler sample’s properties are consistent with each system having the maximum stable angular momentum deficit (AMD; see e.g., Laskar & Petit, 2017) distributed randomly, weighted by mass, between the planets. The sweeping resonances described here conserve angular momentum (and therefore AMD), meaning that these resonances do not produce a systematic signal contrary to the maximum-AMD model. However, as can been seen in Figures 3 and 7, these resonances cause the outer two planets to align and circularize, dumping much of their AMD into the (usually less-massive) inner planet to make its orbit misaligned and eccentric. If the initial distribution of AMD between planets is as assumed by He et al. (2020), where the more massive planets hold more of the system’s AMD, then observations of lower-mass inner planets holding outsized proportions of AMD could be an indication that these sweeping resonances have occurred.
@@
Figure 7: Relation between initial mutual inclination between the two outermost planets and changes in alignment between planets over the course of the simulation for nine systems. The top six systems undergo resonance-mediated misalignments. The bottom three do not. In red circles are, for each simulation, the change in alignment between the inner and middle planets, ($\Delta i_{mut,01}$)–the average mutual inclination at late times between the inner and middle planet minus the initial mutual inclination between the inner and middle planet (see Eq 18). The red dotted line represents the minimum $\Delta i_{mut,01}$ such that the inner planet fails to transit at the end of the simulation (under the simplifying assumption that the middle planet has a transit impact parameter $b_{1}=0$). In blue triangles are, for each simulation, the change in alignment between the middle and outer planets, ($\Delta i_{mut,12}$)–the average mutual inclination at late times between the middle and outer planet minus the initial mutual inclination between the middle and outer planet. The blue dotted line represents the maximum value of $\Delta i_{mut,12}$ as a function of $i_{mut,12,init}$ such that the outer two planets are guaranteed to co-transit at the end of the simulation if $b_{1}=0$. Lines of best fit are plotted for both the red circles and blue triangles. The shaded regions represent the range of $i_{mut,12,init}$ such that all three planets are guaranteed to transit (again, if $b_{1}=0$). The color of the shaded region indicates which planet escapes transit first as $i_{mut,12,init}$ increases–red for the inner planet and blue for the outer planet. The scheme by which orbital parameters were varied between simulations is described in Table 2.
@@ 4.2 The Impact of Static vs Sweeping Resonances
Static resonances notably do not produce the misalignment from evolving $J_{2}$ that sweeping resonances do. For example, the system TOI-2076 (observed by e.g., Osborn et al., 2022; Hedges et al., 2021) lies on such a static resonance as depicted in the bottom middle panel of Figure 6 . In Figure 9, we show the semimajor axis parameter space of TOI-2076 for $J_{2}=10^{-3}$ (left panel) and $J_{2}=0$ (right panel). The location of TOI-2076’s observed semimajor axis ratios is marked with a red dot. In both panels, TOI-2076 lies on a secular resonance with the same expected inner planet maximum inclination. This indicates that there is no expected change to the system as $J_{2}$ evolves from $\sim 10^{-3}$ to 0. Therefore, we do not expect the inner planet to systematically and/or permanently misalign from its companions any more than is provided in its initial conditions. Performing the eigendecomposition described in Section 2, calculating the amplitudes of each of the planets in each of the eigenfrequencies of the $B$ matrix, one finds that all three planets have their maximum amplitude in the same mode for any of the values of $J_{2}$ taken on in Figure 2. This mode’s period can be seen to vary from $10^{4}$ years to $10^{7}$ years in the top row of Figure 8. Therefore, the three planets have their $\Omega$ values precess together forever, with all the angles $\Omega_{1}-\Omega_{0}$, $\Omega_{2}-\Omega_{0}$, and $\Omega_{2}-\Omega_{1}$ librating.
Indeed, in Figure 8, we show the time evolution of one of the simulations of TOI-2076 from the Monte Carlo. In the upper right panel, we can see that the longitudes of ascending nodes between the outer two planets are locked together; however, the upper left panel shows that mutual inclinations do not significantly change. The planets do not misalign. In the bottom middle panel of Figure 7, we can see that even for large initial values of $i_{mut,12}$, no permanent misalignment occurs between the planets in this system–they retain whatever mutual inclinations they are initialized with.
@@
Overplotted on all figures are the Laplace-Lagrange eigenfrequencies $f_{i}$ and $g_{i}$.